# Dynamically Augmented CVaR for MDPs

### v3

@@ 4 Dynamically Augmented CVaR
Let us consider the DRMDP. Theorem 3.3 demonstrates time inconsistency of the static CVaR for MDPs because, for a policy on which CVaR is achieved, current decisions of Nature may depend on future decisions of the DM. The following definition addresses this problem.
@@
where is the set of optimal policies for Nature.
Since this definition also applies to policies According to Theorem 3.1, for each tail risk level there is an optimal nonrandomized policy minimizing the CVaR for the MDP. If the DM plays a nonrandomized risk-independent policy then, in view of Theorem 3.3,
Thus, in addition to being a more natural objective function than CVaR, DCVaR can be used for establishing performance guarantees for CVaR. In addition,
@@
where the equality follows from the existence and definition of an optimal policy for the DM in the DRMDP, and the inequalities follow from the definition of the DCVaR and from the first equality in Theorem A.3. Thus all inequalities in the last formula hold in the form of equalities.
@@
We recall that the function is given, the functions are bounded and continuous in and uniformly in as Values can be computed by value iterations, and the interval can be easily discretized for computations. In addition is the unique bounded solution of (4.2) with
Here we would like to mention that time-inconsistency of the static CVaR for MDPs is well-known, and another objective, called the nested CVaR, is broadly used [29, 30, 31, 33, 34, 35, 36]. One of the definitions of the nested CVaR is relevant to abstract dynamic programming [3] dealing with equations for where is an objective function for a finite or infinite time horizon is the state, is the action, and is a terminal cost. Since each action a defines transition probabilities, can be viewed as a risk measure with respect to this probability. For the nested CVaR this function is the CVaR with respect to a transition probability of future costs depending on the current cost and the evaluation of future total costs. For an MDP, the optimality equation for the nested CVaR for the risk level is
@@
Let us define Then (4.2) becomes
This is the optimality equation for the DRMDP1 defined in the next paragraph. It is easier to deal with the DRMDP1 than with the DRMDP because the functions in concave when is concave in ; Lemma 7.1.
The DRMDP1 is an RMDP defined by the tuple where the state space the action space uncertainty space sets of available actions and uncertainty sets are the same as in RMDP, one-step costs and the transition probability where and Let be the expected total cost for the DRMD1. The following theorem implies that the DRMDP and DRMDP1 represent the same problem.
@@
The following corollary follows directly from Theorem 4.2.
@@
A nonrandomized risk-independent policy for the DM minimizes DCVaR, for a finite or infinite horizon initial state and initial risk level , that is if and only if for an optimal policy for DRMDP1.
@@
Let be a nonrandomized risk-independent policy for the DM, and If for every optimal policy for the DRMDP, then
@@
The sets of optimal actions for the DM are
@@
which is the minimax equation for the sequential deterministic game in which the DM chooses actions, and Nature chooses a transition from the set of transitions having positive probability. The DM is trying to maximize a feasible path while Nature is trying to minimize it. For a horizon a path is the sequence and for it is For a given initial state a path is called feasible if and for all integer values The lengths of finite and infinite-horizon paths are defined by formulae (2.6) and (2.7) for finite and infinite horizon problems respectively. This is a special case of an RMDP considered in Appendix A. In this case, the state space is finite, action sets for the DM are uncertainty sets for Nature are one step costs are and all moves are deterministic to the next states selected by Nature. This problem can be solved easily by value iteration.
We notice that, if then is the expected total discounted reward for the policy since if
@@ 5 Formulation of the Main Result
Let us consider the DRMDP1. We recall that where the function is continuous in When everywhere in the rest of this paper, except Appendix A, we assume that Assumption 5.1 holds. For example, this assumption holds when the function is nonincreasing in and concave in for all In particular, this assumption holds if the final payoff does not depend on
@@
In this section we introduce Algorithm DCVaR constructing a nonrandomized policy minimizing for and The case is addressed in the previous section. In addition, This, in view of (4.1), minimizes DCVaR within the set of all policies for the DM. Thus, let
@@
derivatives respectively, where the concavity of implies If is a real-valued continuous concave function defined on the interval we set and For a function of multiple variables, say with all variables except one, say , we shall consider a function in for fixed values of all other parameters. The right and left derivatives in if they exist, will be also denoted as and respectively. If a derivative in exists for some , and it is denoted by which means that We usually apply these notations to functions of two variables where and We recall that, for a real-valued continuous concave function defined on an interval, the right and left derivatives always exist, the function is right-continuous and lower semicontinuous, and the function is left-continuous and upper semicontinuous. A concave function on an interval is differentiable everywhere except for at most a countable set. We also denote by the superdifferential of in at the point In particular, if exists, then We shall also apply this notation for the superdifferential of functions of one and two variables.
@@
Let us consider the sets of optimal actions defined in (4.6), where If for every optimal policy of Nature, at each finite epoch such that a policy of the DM chooses actions from the state then this policy minimizes in
The following equalities hold in view of Hiriart-Urruty and Lemaréchal [17, p.28]:
@@
Algorithm DCVaR for running an optimal policy for an MDP with a finite state space action sets available at states transition probability costs transition probability and discount factor
Inputs: Initial state , tail risk level the horizon length or the value functions if or the value function if After an action is selected at the state where is a nonnegative integer such that the next state such that becomes known.
@@ 2.2. Compute:
2.4. If then stop; otherwise, set and choose an arbitrary
If then the algorithm stops after N iterations and returns a finite sentence of actions implementing an optimal policy at the steps If then the algorithm returns an infinite sequence of actions implementing an optimal policy at the steps and the algorithm can be stopped under a finite number of iterations because the impact of additional steps will be negligibly small due to the discount factor and estimations of stopping times are standard.
As an illustration of Algorithm CVaR, Figure 1 shows the two main cases addressed by the algorithm. Figure (1(a)) is an example of a case where there is a unique tail risk level being identified in step 2.3, and Figure (1(b)) corresponds to the other case where there is an interval with the desired slope value.
@@ 2.4. If then stop; otherwise, set and choose an arbitrary
Tail risk levels assigned by Nature are not available to the Decision Maker when However, by using formula (5.3), which is based on the analysis of optimal policies by Nature, at step 2.3 the algorithm detects either the tail risk level or an interval, in which is located, and each internal point of this interval defines the set of optimal actions for the DM, which are also optimal if Nature selects another tail risk level from this interval. The main result of this paper is summarized in the following theorem, whose proof is provided in Section 8.
@@
For or and Algorithm DCVaR generates a nonrandomized risk-independent policy minimizing DCVaR for which
@@ 6 Properties of the Mass Transfer Problems Solved by Nature
This section studies the problem Nature solves at each horizon or The section studies problem (6.1) motivated by the maximization operation in equation (4.4). The results of this section are self-contained. Continuity and concavity in of the function in formula (6.1) are assumed because, as shown in the next section, these assumptions are satisfied by the functions in (4.5). This theorem is the key fact on which formula (5.3) in Algorithm DCVaR is based.
A function where is called piecewise linear, if there is a finite sequence of increasing numbers with and such that the function is linear on each interval
Let be a real-valued function such that for each fixed the function is continuous and concave in For a probability distribution on that is, for all and , let
@@
A function is called feasible for problem (6.1) if for each the function is Borel-measurable on and for all For example, is a feasible function, and, if then is an optimal solutions at to the problem defined in (6.1). A feasible function for problem (6.1) is called a solution to problem (6.1) if for all
@@
We observe that is a continuous compact-valued set-valued mapping of into the set of subset of In addition for all and the function is continuous on Therefore, the Berge maximum theorem implies that the maximum in (6.1) exists, the function is continuous, and the set-valued mapping is upper-semicontinuous at all In view of the Kuratowsli-Ryll-Nardzewski measurable selection theorem, there is a measurable optimal solution b(y), which means that all functions are measurable in Let and where Then Then where the second inequality follows from concavity of ∎
We notice that concavity of implies continuity of but we did not use this fact in the proof of Theorem 6.1. The following theorem describes the properties of the solutions to problem (6.1).
@@
Problem (6.1) can be simplified to a problem which has a natural interpretation. First, without loss of generality we can assume that as this takes place for the function Indeed, we can always consider the objective function Then all the objective and value functions will be shifted by the constant and the solution sets won’t change. Second, we assume for all without loss of generality because the points with can be excluded from
@@
Problem (6.7) describes the following optimal mass transfer problem. For let us consider intervals in which we call sources. These sources can be interpreted as cylindrical vessels of the same diameter filled with a liquid up to the height Let There is an additional interval which we call the destination. This interval can be interpreted as an empty vertical cylindrical vessel with the same diameter and with a height 1. The total value of the liquid at source from level 0 up to the level is where is a concave function with The goal is to fill up the destination by the liquid from sources to maximize for each the total value of the liquid in the destination from the level 0 up to the level . Any amount of the liquid can be taken from any part of each source.
@@
and the set is countable. Thus, both functions and where describe the nonlinear unit cost of the liquid at the source depending on the height These unit costs may be have arbitrary signs (positive, negative, or 0), but they are nondecreasing functions of the hight since the function is convex in
Let us prove (6.2) and (6.5). If then these formulae hold in the form Let Since the function is convex, the most valuable part of the liquid of the volume at the source is the interval whose value is Thus, if the amounts of liquid should be taken by an optimal policy for the level from sources then the decision to take all the liquid from the interval from each source is optimal, and almost everywhere for Since the function is concave and therefore the function is left-continuous, this inequality holds for all In particular, for So. (6.2) is proved.
Equality (6.2) can be written as for In particular,
and the strict inequality is impossible. Indeed, if the strict inequality holds, then there exists such that for all
Since there exists such that Let Then we consider the feasible solution with the same decisions outside of the interval of and allocated all possible resource from the source when the level is in this interval. The formal definition is
Then
Let us consider the function where Then for we have For we have for and which implies and therefore Thus,
which is impossible. Thus, equation (6.9) and therefore equation (6.5) are proved.
The proof of (6.6) is similar. In particular, for all is the only solution for and (6.6) holds in the form of So, we consider The same arguments as for (6.2) imply (6.3). Therefore, where is a solution of (6.7) corresponding to a solution of (6.1). The similar argument as in the proof of (6.5) imply that the strict inequality is impossible because otherwise there is a solution such that for some Thus, (6.10) and therefore (6.6) hold. Equations (6.5) and (6.6) imply (6.4). The last claim is correct because if functions are piecewise linear in for each , then formulae (6.5) and (6.6) imply that the concave functions and take finite numbers of values when therefore, the concave function is piecewise linear on the interval ∎
Let for all
@@
It is possible to construct all solutions of problem (6.1). We do not use these solutions in this paper. So, we we do not provide detailed proofs. We recall that it does not matter which solution is chosen when and is the only solution for Let Then, let us consider two cases. In case 1 does not belongs to an interval on which the concave function is linear. In case 2 belongs to an interval on which the function is linear. In case 2 we choose in the way that this is the maximal open interval containing on which is linear. In case 1 there is a unique solution at . In case 2, where is the set of all states such that there are points such that and the function is linear on with its derivative on this interval equal to either or to In this case we consider being the maximal interval satisfying this condition. Of course, if In case 2, and is a solution at iff: (i) for and (ii) and for We notice that, in case 2, if is a singleton, then the solution of problem (6.1) at is unique.
@@
In order for solutions to have physical meaning of moving the liquid from sources to the destination, the functions or should be nondecreasing in To achieve this, it is sufficient to set for all and to define proprietary the values for the situations when the function belongs to an interval on which the function is linear and is not a singleton. In this case, for example, we can move all the liquid from sources sequentially. For example, sources with smaller numbers can be used first. Such solutions are nondecreasing in Of course, there are other ways to construct nondecreasing solutions.
@@
1. Merge sequences into a single finite sequence By doing this, merge the intervals with the same slopes. This means that any final set of pairs with distinct indexes where and should be replaced with the single pair

### v4

@@ 4 Dynamically Augmented CVaR
Let us consider the DRMDP with if and if Theorem 3.3 demonstrates time inconsistency of the static CVaR for MDPs because, for a policy on which CVaR is achieved, current decisions of Nature may depend on future decisions of the DM. The following definition addresses this problem. Optimal decisions of Nature do not depend on future actions of the DM. Therefore, it is natural for Nature to make optimal decisions.
@@
where is the set of optimal policies for Nature for the -horizon problem.
Since this definitions applies to policies According to Theorem 3.1, for each tail risk level there is an optimal nonrandomized policy minimizing the CVaR for the MDP. If the DM plays a nonrandomized risk-independent policy then, in view of and Theorem 3.3,
Thus, in addition to being a more natural objective function than CVaR, DCVaR can be used for establishing performance guarantees for CVaR. In addition, for the DRMDP
@@
where the equality follows from the existence and definition of an optimal policy for the DM in the DRMDP, and the inequalities follow from the definition of the DCVaR and from the first equality in Theorem A.7. Thus all inequalities in (4.2) hold in the form of equalities.
@@
We recall that the bounded continuous function is given, the functions are bounded and continuous in and uniformly in as see (A.4). Values can be computed by value iterations, and the interval can be easily discretized for computations. In addition is the unique bounded solution of (4.3) with
Here we would like to mention that time-inconsistency of the static CVaR for MDPs is well-known, and another objective, called the nested CVaR, is broadly used [33, 34, 35, 37, 38, 39, 40]. One of the definitions of the nested CVaR is relevant to abstract dynamic programming [3] dealing with equations for where is an objective function for a finite or infinite time horizon is the state, is the action, and is a terminal cost. For the nested CVaR, the function is CVaR of with respect to a transition probability where is the one-step cost. For an MDP, the optimality equation for the nested CVaR for the risk level is
@@
Let us define for or and Then (4.3) becomes
This is the optimality equation for the DRMDP1 defined in the next paragraph. It is easier to deal with the DRMDP1 than with the DRMDP because, as stated in Lemma 7.1, the value functions are concave in when is concave in as assumed in Assumption 5.1.
Let us introduce the DRMDP1 as an RMDP defined by the tuple where the state space the action space uncertainty space sets of available actions and uncertainty sets are the same as in the DRMDP, one-step costs and the transition probability where and Let be the expected total cost for the DRMD1. The following theorem implies that the DRMDP and DRMDP1 correspond to the same problem when where
@@
The following two corollaries follow directly from Theorem 4.2 and from formula (4.2), which holds in the form of equalities.
@@
For a finite or infinite horizon initial state and initial risk level , the equality holds for a nonrandomized risk-independent policy for the DM if and only if for every optimal policy for Nature.
@@
For a finite or infinite horizon let be a nonrandomized risk-independent policy for the DM, and If for every optimal policy for Nature, then
and, in particular, is DM’s optimal policy for the initial state
@@
According to (A.5) and (A.6), the sets of optimal actions for the DM are
@@
The following theorem provides a sufficient condition under which a policy for the MDP minimizes DCVaR. The assumption that the policy is nonrandomized is not needed. We formulate this assumption only because this theorem is applied to a nonrandomized policies in this paper.
Theorem 4.5.
For a nonrandomized risk-independent policy initial state tail risk level and time horizon or if for every persistently optimal policy for Nature and for every nonnegative integer then formulae (4.6) hold.
Proof.
@@
which is the minimax equation for the sequential deterministic game in which the DM chooses actions, and Nature chooses a transition from the set of transitions having positive probability. The DM is trying to minimize the length of a feasible path while Nature is trying to maximize it. For a horizon a path is a sequence and for it is For a given initial state a path is called feasible if and for all integer values The lengths of finite and infinite-horizon paths are defined by formulae (2.2) and (2.3) for finite and infinite horizon problems respectively. This is a special case of an RMDP considered in Appendix A. In this case, the state space is finite, action sets for the DM are uncertainty sets for Nature are one step costs are and all moves are deterministic to the next states selected by Nature. This problem can be solved easily by value iteration.
We notice that, if then all uncertainty sets can be reduced to singletons where such that if Thus, if then we deal with an MDP with expected total discounted costs.
@@ 5 Formulation of the Main Result
Let us consider the DRMDP1. We recall that where the function is continuous in When everywhere in the rest of this paper, except Appendix A, we assume that Assumption 5.1 holds. For example, this assumption holds when the function is nonincreasing in and concave in for all In particular, this assumption holds if the final payoff does not depend on that is,
@@
In this section we introduce Algorithm DCVaR constructing for and a nonrandomized policy minimizing among all policies In view of (4.1), this means that The case is addressed in the previous section. Let
@@
derivatives respectively, where the concavity of implies If is a real-valued continuous concave function defined on the interval we set and For a function of multiple variables, say with all variables except one, say , being discrete, we shall consider a function in for fixed values of all other parameters. The right and left derivatives in if they exist, will be also denoted as and respectively. If a derivative in exists for some , and it is denoted by which means that We usually apply these notations to functions of two variables where and We recall that, for a real-valued continuous concave function defined on an interval, the right and left derivatives always exist, the function is right-continuous and lower semicontinuous, and the function is left-continuous and upper semicontinuous. A concave function on an interval is differentiable everywhere except for at most a countable set. We also denote by the superdifferential of in at the point In particular, if exists, then We shall also apply this notation for the superdifferential of functions of one and two variables.
@@
Let us consider the sets of optimal actions defined in (4.8). According to Theorem 4.5, if for every optimal policy of Nature, at each finite epoch such that a policy of the DM chooses actions from the state then this policy minimizes in
The following equalities hold in view of Hiriart-Urruty and Lemaréchal [18, p.28]:
@@
Algorithm DCVaR for running an optimal policy for an MDP with a finite state space finite action sets available at states transition probability one-step costs and discount factor where if
Inputs: Initial state tail risk level the horizon length or the value functions if or the value function if After an action is selected at the state where is a nonnegative integer such that the next state such that becomes known.
@@ 2.2. Compute:
2.4. If then stop; otherwise, choose an arbitrary optimal action set and then set
If then the algorithm stops after N iterations and returns a finite sentence of actions
implementing an optimal policy at the steps If then the algorithm returns an infinite sequence of actions implementing an optimal policy at the steps and the algorithm can be stopped after a finite number of iterations because the impact of additional steps will be negligibly small due to the discount factor and estimations of stopping times are standard.
As an illustration of Algorithm DCVaR, Figure 1 shows the two main cases addressed by the algorithm. Figure (1(a)) is an example of a case where there is a unique tail risk level being identified in step 2.3, and Figure (1(b)) corresponds to the other case where there is an interval with the desired slope value.
@@ 2.4. If then stop; otherwise, choose an arbitrary optimal action set and then set
Tail risk levels assigned by Nature are not available to the DM when However, by using formula (5.3), which is based on the analysis of Nature’s optimal policies, at step 2.3 the algorithm detects either the tail risk level or an interval, in which is located, and every optimal action for an internal point of this interval is also optimal if Nature selects another tail risk level from this interval. The main result of this paper is summarized in the following theorem, whose proof is provided in Section 8.
@@
For or and Algorithm DCVaR generates a nonrandomized risk-independent policy minimizing DCVaR, that is,
@@ 6 Properties of the Mass Transfer Problems Solved by Nature
This section studies the problem Nature solves at each horizon or In particular, it studies problem (6.1) motivated by the maximization equation in (4.7). The results of this section are self-contained. Continuity and concavity in of the function in formula (6.1) are assumed because, as shown in the next section, these assumptions are satisfied by the functions and from (4.7). Tneorem 6.2 is the key fact on which formula (5.3) in Algorithm DCVaR is based.
A function where is called piecewise linear, if there is a finite sequence of increasing numbers with and such that the function is affine on each interval
Let be a real-valued function such that for each fixed this function is continuous and concave in For a probability distribution on that is, for all and , let
@@
A function is called feasible for problem (6.1) if for each the function is Borel-measurable on and for all For example, is a feasible function. A feasible function for problem (6.1) is called a solution to problem (6.1) if for all
@@
Without loss of generality we assume that for all because, if then state can be excluded from the set We observe that is a continuous compact-valued set-valued mapping of into the set of subsets of In addition for all and the function is continuous on Therefore, the Berge maximum theorem implies that are nonempty compact sets for all the function is continuous, and the set-valued mapping is upper-semicontinuous at all In view of the Kuratowsli-Ryll-Nardzewski measurable selection theorem, there is a Borel-measurable optimal solution which implies that all functions are Borel-measurable in Let and where Then Then where the second inequality follows from concavity of ∎
We notice that concavity of implies continuity of but this fact is not used in the proof of Theorem 6.1. The following theorem describes the properties of the solutions to problem (6.1).
@@
Problem (6.1) can be simplified to a problem which has a natural interpretation. First, without loss of generality we can assume that Indeed, we can always consider the objective function Then all the objective and value functions will be shifted by the constant and the solution sets won’t change. Second, in the rest of this section we assume without loss of generality that for all
@@
Problem (6.7) describes the following optimal mass transfer problem. For let us consider intervals in which we call sources. These sources can be interpreted as vertical cylindrical vessels of the same diameter filled with a liquid up to heights Let There is an additional interval which we call the destination. This interval can be interpreted as an empty vertical cylindrical vessel with the same diameter and with height 1. The total value of the liquid at source from level 0 up to the level is where is a concave function with The goal is to fill up the destination by the liquid from sources to maximize for each the total value of the liquid in the destination from the level 0 up to the level . Any amount of the liquid can be taken from any part of each source.
@@
and the set is countable. Thus, both functions and where describe the nonlinear unit cost of the liquid at the source depending on the height These unit costs may have arbitrary signs (positive, negative, or 0), but they are nondecreasing functions of the hight since the function is concave in
Let us prove (6.2) and (6.5). If then these formulae hold in the form Let Since the function is concave, the most valuable part of the liquid of the volume at the source is the interval whose value is Thus, if the amounts of liquid should be taken by an optimal policy for the level from sources then the decision to take all the liquid from the interval from each source is optimal, and for all In particular, for So. (6.2) is proved.
Inequality (6.2) can be written as for if In particular, and the strict inequality is impossible. Indeed, if the strict inequality holds, then concavity of and imply the existence of such that for all Since there exists such that Let Then we consider the feasible solution with the same decisions outside of the interval of and allocating all possible resource from the source when the level is in this interval. The formal definition is: for
Let us verify that is a feasible function for problem (6.7). We need to verify the properties of for Let Then In addition, Thus, is feasible, and
which is impossible since is an optimal value, where the last equality in (LABEL:eq6fin) follows from
The proof of (6.6) is similar. In particular, for all is the only solution for and (6.6) holds in the form of So, we consider The same arguments as for (6.2) imply (6.3). Therefore, where is a solution of (6.7) corresponding to a solution of (6.1). The similar argument as in the proof of (6.5) imply that the strict inequality is impossible because otherwise there is a solution such that for some Thus, (6.10) and therefore (6.6) hold. Equations (6.5) and (6.6) imply (6.4). The last claim is correct because, if functions are piecewise linear in for each , then formulae (6.5) and (6.6) imply that the functions and take finite numbers of values when Therefore, the concave function is piecewise linear on the interval ∎
@@
It is possible to construct all solutions of problem (6.1). We do not use these solutions in this paper. So, we do not provide details on constructing them. If then is the unique solution of problem (6.7). If then is the unique solution of (6.7). Thus, is a solution of problem (6.1) for and for this solution is unique. Let Then, let us consider two cases. In case 1, does not belongs to an interval on which the concave function is linear. In case 2, belongs to an interval on which the function is linear. In case 2 we choose in the way that this is the maximal open interval containing on which is linear. In case 1, there is a unique solution at . In case 2, where is the set of all states such that there are points such that and the function is linear on with its derivative on this interval equal to In this case we consider being the maximal open interval satisfying this condition. If then In case 2, and is a solution at iff: (i) for and (ii) and for We notice that, in case 2, if is a singleton, then if and is the unique solution of problem (6.1) at
@@
In order for solutions to have physical meaning of moving the liquid from sources to the destination, the optimal functions should be nondecreasing in To achieve this, it is sufficient to define appropriately the values for when belongs to an interval on which the function is linear, and is not a singleton. In this case, for example, we can move the liquid from sources to the destination sequentially. Sources with smaller numbers can be used first. Such solutions are nondecreasing in Of course, there are other ways to construct nondecreasing solutions.
@@
1. Merge sequences into a single finite sequence By doing this, merge the intervals with the same slopes. This means that any final set of pairs with distinct indexes and the same slope where and should be replaced with the single pair